OpenStudy (anonymous):

The diagonal of a rectangle is 10. What is the area of the rectangle? A. 24 B. 48 C. 50 D. 100 E. cannot be determined from the info given

OpenStudy (mandre):

E. You need one more side and angle.

OpenStudy (mandre):

or another side

OpenStudy (dumbcow):

 Area = xy x^2 + y^2 = 10^2 the possible sides form a circle with radius 10

OpenStudy (mathstudent55):

These three rectangles have congruent diagonals. As you can see their areas are quite different. The conclusion is that just knowing the length of the diagonal of a rectangle is not enough information to find its area. 

OpenStudy (mathmate):

Note: If the sides have to be integers, then there is a unique answer.

OpenStudy (jdoe0001):

if the diagonal is 10 we know the other 2 sides will be less than 10, that is a < 10, b < 10 if you pick a value for either, you can get the value of the 2nd one by using pythagorean theorem and regardless of what value you pick, so long is less than 10 the Area of the rectangle will remain the same

OpenStudy (anonymous):

So I can't use the 30-60-90 triangle formula?

OpenStudy (dumbcow):

its kind of a trick question...by itself there is not enough info to know the Area but since they give possible Areas, you can determine which possible Area is correct example: A = 48 8*6 = 48 and 8^2 + 6^2 = 100

OpenStudy (jdoe0001):

say if I were to say pick for my side b = 5, my c = 10 that usually means a triangle of 30-60-90, meaning my \(\bf a = 5\sqrt{3}\) using that I can find the area of that triangle, double that, and you get the area of the rectangle

OpenStudy (jdoe0001):

the Area will be the same, no matter at what angle you arrange the diagonal, you see

OpenStudy (jdoe0001):

because the diagonal carries a ratio for the other 2 sides

OpenStudy (anonymous):

Ah I get it now! Thank you!

OpenStudy (phi):

unless they tell you the sides are integer length, there is not enough info to know the area of the rectangle.

OpenStudy (dumbcow):

@jdoe, not to confuse the situation but by changing the angle (shape of rectangle) the Area does definitely changes

OpenStudy (jdoe0001):

yes it does, ok, my bad, was triple checking